OpenStudy (anonymous):

how to find solution set of the equation:3^2+x=5^x?

OpenStudy (john_es):

Do you mean \[3^2+x=5^x\Rightarrow 9+x=5^x?\]

OpenStudy (anonymous):

@John_ES no, 2+x is actuallly exponent of \[3^2+^x=5^x\]

OpenStudy (john_es):

\[\log(3^{2+x})=\log(5^x)\Rightarrow (2+x)\log3=x\log5\Rightarrow x=\frac{2\log3}{\log5-\log3}\] \[x=\frac{2\log3}{\log(5/3)}\approx4.3\] The logs are decimal logarithms..